A frog jumps and takes 2 seconds to land. The path of the frog’s flight is best modeled by the graph of f(x) = -x2 + 2x. What is the maximum height of the jump?


1
Expert's answer
2022-04-21T11:49:26-0400

"f(x)=-x^2+2x=-(x-1)^2+1."

"f_{max}=f(1)=1."